OpenStudy (anonymous):

7x-2y=42. Can someone help explain thoroughly how to find the x and y intercepts in problems like this?

OpenStudy (anonymous):

Maybe this link can help you http://cs.selu.edu/~rbyrd/math/intercept/

OpenStudy (lgbasallote):

i can give you an example... suppose you have 5x + 4y = 20 to find x-intercept let y = 0 then solve for x so when y = 0 5x + 4(0) = 20 5x = 20 x = 20/5 x = 4 so x-intercept is (4,0) to find y-intercept let x = 0 when x = 0 5(0) + 4y = 20 4y = 20 y = 20/4 y = 5 so y-intercept is (0,5) does that help?

OpenStudy (anonymous):

If i did it like that it would be 7x - 2(0) = 42 X = 6. And 7(0) - 2y = 42 would be y = - 21. So it would be: (6,0) and (0,-21)

OpenStudy (lgbasallote):

that's right!

OpenStudy (anonymous):

Thanks. :)

OpenStudy (lgbasallote):

welcome